Fermi National Laboratory
 

Operation Summary for March 10, 2010 - March 12, 2010

 

Wednesday March 10

The day shift began with Operations monitoring the Tevatron (TeV) store 7663 (262.48E30) and the Recycler stash (64.2E10), with the Antiproton Source stacking (9.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, a Booster RF station (BRF11) did not survive the fluctuation in LCW flow experienced on the previous shift. Operators disconnected it from the anode supply; RF personnel repaired and reconnected the station in the afternoon. EE Support personnel investigated and addressed problems with an MI extraction kicker (KPS62A).

During the evening shift, Operations attributed a minimal amount of downtime to a faulty Pbar lithium lens interlock sensor and various Booster and Linac RF station trips.

Thursday March 11

The midnight shift began with Operations monitoring store 7663 (126.16E30) and the stash (218.1E10), with Pbar stacking (17.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7663 at 6:30 AM. Operations reported that the TeV would remain at 150 GeV for the duration of the CUB chiller startup. Operators prepared the Pbar Rings for access and keys were issued at 7 AM. CUB chiller maintenance began shortly after 7 AM.

During the day shift, chiller #2 was brought online by 9:00 AM and the Pbar work was done by 10:00 AM. A short access was made into Tevatron A-E to investigate a high water alarm. The sump pump was ground faulted, so a temporary pump was added until further repairs can be made. Stacking was established by 10:15 AM and TeV startup was underway by 1:15 PM. Beam was delivered for stacking and to users when available.

During the evening shift, shot setup was underway by 4:15 PM. Store 7667 was colliding with an AIL=351E30 by 5:30 PM. There was minor downtime due to various RF trips in Linac and Booster. Beam was delivered for stacking and to users when available.

Friday March 12

The midnight shift began with Operations monitoring store 7667 (135.13E30) and the stash (343.1E10), with Pbar stacking (17.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, store 7667 was terminated as scheduled at 3:45 AM and shot setup followed. Store 7668 was colliding with an AIL=347E30 by 5:00 AM. Booster RF station 15 (BRF15) modulator was having problems and was disconnected from the anode. Beam was delivered to the Pbar Source, NuMI, MiniBooNe, and MTest when it was available. Pbar transfers to the Recycler were made when appropriate.

The Plans for Friday and the Weekend

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/12/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory